# Single-photon scattering in giant-atom topological-waveguide-QED systems

### v1

@@ III.3.1 The AAAA-coupling configuration
Figure 5: The reflection coefficient $R_{2}$ for the AAAA-coupling configuration as a function of $\Delta_{k}$ and $\Delta$. The dimerization parameter is (a) $|\delta|=0.5$. The profiles of panel (a) at several values of $\Delta$ are shown in (b). Other parameters used are $g=0.01J$ and $d=2$.
For the AAAA-coupling configuration, we have $\mu_{i}=1$ and $\nu_{i}=0$ (for $i=1-4$) in Eqs. (35a)-(36f). Figure 5(a) shows the reflection coefficient as a function of $\Delta_{k}$ and $\Delta$. Similar to the AA-coupling configuration in the single-giant-atom case, the sign of $\delta$ has no influence on the single-photon scattering behavior for the AAAA-coupling configuration. Thus, we only need to analyze the influence of $|\delta|$ on the scattering spectra. However, the reflection spectra show more abundant line shapes due to the collective effect of the two giant atoms.
In Fig. 5(b), we plot the profile of Fig. 5(a) at several frequency detunings in the region of $\Delta\in(J,2J)$, we find that when $\Delta\approx 1.20J$ (the blue dashed curve), the characteristic quantities in Eqs. (37a)-(37e) become $\Delta_{\text{S}L,k_{\Delta}}=\Delta_{\text{A}L,k_{\Delta}}=\Gamma_{e}$, $\Gamma_{\text{S},k_{\Delta}}=0$, $\Gamma_{\text{A},k_{\Delta}}=4\Gamma_{e}$, and $g_{\text{SA},k_{\Delta}}=\Gamma_{\text{SA},k_{\Delta}}=0$. Therefore, the symmetric (antisymmetric) state is decoupled from (coupled to) the waveguide. This feature results in the appearance of the Lorentzian line shape centered at $\Delta_{k}=\Gamma_{e}$ with the linewidth of $4\Gamma_{e}$. When $\Delta\approx 1.58J$ (the purple dot-dashed curve), the six characteristic quantities are $\Delta_{\text{S}L,k_{\Delta}}=\Delta_{\text{A}L,k_{\Delta}}=\Gamma_{\text{S},k_{\Delta}}=\Gamma_{\text{A},k_{\Delta}}=g_{\text{SA},k_{\Delta}}=\Gamma_{\text{SA},k_{\Delta}}=0$. In this case, both the symmetric and antisymmetric states are decoupled from the SSH waveguide. Therefore, the incident photon in the waveguide can be completely transmitted. When $\Delta=1.15J$ (the red solid curve) and $1.49J$ (the black dotted curve), the spectra become asymmetric, with a minimal reflection located at $\Delta_{k}=\Gamma_{e}\{\sin(k_{\Delta}d)-[1+\cos(k_{\Delta}d)]\tan(2k_{\Delta}d)\}$.
In the following analyses, we can show that in some regions the reflection spectrum around the minimal reflection can be approximated as a standard Fano line shape Fano1961 ; Miroshnichenko2010 . To this end, we decompose the reflection amplitude in Eq. (LABEL:r2) as $r_{2,\pm}^{\text{AAAA}}=r_{2,+}^{\text{AAAA}}+r_{2,-}^{\text{AAAA}}$, where
@@
Figure 6: The reflection coefficient $R_{2}$ for the ABAB-coupling configuration as a function of $\Delta_{k}$ and $\Delta$. The dimerization parameter is (a) $\delta=0.5$ and (b) $\delta=-0.5$. The profiles of panels (a) and (b) at several values of $\Delta$ are shown by the curves in (c) and (d). Other parameters used are $g=0.01J$ and $d=2$.
@@ III.3.2 The ABAB-coupling configuration
For the ABAB-coupling configuration, the dimensionless parameters become $\mu_{1(3)}=\nu_{2(4)}=1$ and $\nu_{1(3)}=\mu_{2(4)}=0$ in Eqs. (35a)-(36f). To see the influence of $\Delta$ and $\delta$ on the single-photon scattering process, in Figs. 6(a) and 6(b) we show the reflection coefficient as a function of $\Delta_{k}$ and $\Delta$. It can be seen that the line shapes of the reflection spectra are sensitive to the sign of $\delta$, which is different from the AAAA-coupling configuration. To show the details more clearly, in Figs. 6(c) and 6(d), we plot the profiles of Figs. 6(a) and 6(b) at different frequency detunings in the region $\Delta\in(J,2J)$, respectively.
@@
We can see from the red solid curve in Fig. 6(c) [Fig. 6(d)] that, when $\delta=0.5$ ($-0.5$) and $\Delta\approx 1.20J$, the characteristic quantities are $\Delta_{\text{S}L,k_{\Delta}}=\Delta_{\text{A}L,k_{\Delta}}\approx-0.96\Gamma_{e}$ ($0.47\Gamma_{e}$), $\Gamma_{\text{S},k_{\Delta}}\approx 0$, $\Gamma_{\text{A}k_{\Delta}}\approx 2.82\Gamma_{e}$ ($0.46\Gamma_{e}$), and $g_{\text{SA},k_{\Delta}}=\Gamma_{\text{SA},k_{\Delta}}=0$. Therefore, the symmetric (antisymmetric) state is decoupled from (coupled to) the SSH waveguide. This makes the reflection spectrum exhibit the Lorentzian line shape centered at $\Delta_{k}=-0.96\Gamma_{e}$ ($0.47\Gamma_{e}$) with the linewidth of $2.82\Gamma_{e}$ ($0.46\Gamma_{e}$), as shown by the red solid curve. However, when $\delta=0.5$ ($-0.5$) and $\Delta\approx 1.58J$, i.e., the purple dot-dashed curve, the characteristic quantities become $\Delta_{\text{S}L,k_{\Delta}}=\Delta_{\text{A}L,k_{\Delta}}\approx-0.32\Gamma_{e}$ ($-0.95\Gamma_{e}$), $\Gamma_{\text{S},k_{\Delta}}\approx 7.79\Gamma_{e}$ ($5.26\Gamma_{e}$), $\Gamma_{\text{A}k_{\Delta}}\approx 0$, and $g_{\text{SA},k_{\Delta}}=\Gamma_{\text{SA},k_{\Delta}}=0$. In this case, the antisymmetric (symmetric) state is decouped from (coupled to) the SSH waveguide, and then the reflection spectrum also exhibits the Lorentzian line shape but centered at $\Delta_{k}=-0.32\Gamma_{e}$ ($-0.95\Gamma_{e}$) with the linewidth of $7.79\Gamma_{e}$ ($5.26\Gamma_{e}$), as shown by the purple dot-dashed curve. When $\delta=\pm 0.5$ and $\Delta=1.49J$ (the black dotted curves), the spectra around the minimal reflection can be approximated as the asymmetric Fano line shapes. Here, the reflection spectrum takes the minima value ($R_{2}=0$) at $\Delta_{k}=\Gamma_{e}\{\sin(k_{\Delta}d-\phi_{k_{\Delta}})-[1+\cos(k_{\Delta}d-\phi_{k_{\Delta}})]\tan(2k_{\Delta}d)\}$. Specially, when $\delta=\pm 0.5$ and $\Delta\approx 1.27J$ (the blue dashed curves), the reflection spectra show the asymmetric Fano line shape and complete transmission, respectively. Similar to the AAAA-coupling configuration, we rewrite the reflection amplitude given by Eq. (LABEL:r2) as the superposition of two Lorentz spectra $r_{2,\pm}^{\text{ABAB}}=r_{2,+}^{\text{ABAB}}+r_{2,-}^{\text{ABAB}}$, where the Lorentz-type amplitudes are given by
@@
Figure 7: The reflection coefficient $R_{2}$ for the AABB-coupling configuration as a function of $\Delta_{k}$ and $\Delta$. The dimerization parameter is (a) $\delta=0.5$ and (b) $\delta=-0.5$. The profiles of panels (a) and (b) at several values of $\Delta$ are shown by the curves in (c) and (d). Other parameters used are $g=0.01J$ and $d=2$.
To obtain the Fano line shapes for the reflection spectra, we need to ensure the condition $\Gamma_{\pm}\gg\Gamma_{\mp}$. Under this condition, the reflection spectrum near $\Delta_{\mp}$ can be approximated as the Fano line shape characterized by Eq. (40). Specifically, when the dimerization parameters $\delta=\pm 0.5$ and $\Delta$ is taken as $\Delta/J\in(1,1.02)\cup(1.48,1.58)\cup(1.58,1.68)\cup(1.99,2)$, the width $\Gamma_{\pm}$ satisfy $\Gamma_{+}\gg\Gamma_{-}$. Thus, the reflection spectra exhibit the Fano line shapes around the reflection minima, as shown by the black dotted curve in Figs. 6(c) and 6(d). When the dimerization parameters $\delta=\pm 0.5$ and $\Delta/J\in(1.11,1.20)\cup(1.20,1.27)\cup(1.27,1.30)\cup(1.82,1.89)\cup(1.89,1.94)$, we have $\Gamma_{-}\gg\Gamma_{+}$, and then the reflection spectrum exhibits the Fano line shape, as shown by the blue dashed curve in Fig. 6(c) (i.e., $\delta=0.5$ and $\Delta\approx 1.27J$). When $\delta=-0.5$ and $\Delta\approx 1.27J$, the characteristic quantities in Eqs. (37a)-(37e) become $\Delta_{\text{S}L,k_{\Delta}}=\Delta_{\text{A}L,k_{\Delta}}=\Gamma_{\text{S},k_{\Delta}}=\Gamma_{\text{A},k_{\Delta}}=g_{\text{SA},k_{\Delta}}=\Gamma_{\text{SA},k_{\Delta}}=0$. Therefore, both the symmetric and antisymmetric states are decoupled from the SSH waveguide. In this case, the input photon can be completely transmitted the waveguide, as shown by the blue dashed curve in Fig. 6(d).
@@ III.3.3 The AABB-coupling configuration
For the AABB-coupling configuration, the dimensionless parameters in Eqs. (35a)-(36f) are $\mu_{1(2)}=\nu_{3(4)}=1$ and $\nu_{1(2)}=\mu_{3(4)}=0$. Figures 7(a) and 7(b) show the reflection coefficient as a function of $\Delta_{k}$ and $\Delta$. We find that the sign of $\delta$ can affect the reflection spectra of the AABB-coupling configuration, but not affect the location of the reflection peak ($R_{2}=1$). This is because the Lamb shifts $\Delta_{L1,k_{\Delta}}=\Delta_{L2,k_{\Delta}}=\Gamma_{e}\sin(k_{\Delta}d)$ are independent of $\delta$. The feature is different from the ABAB-coupling configuration.
Figure 8: The reflection coefficient $R_{2}$ for the ABBA-coupling configuration as a function of $\Delta_{k}$ and $\Delta$. The dimerization parameter is (a) $\delta=0.5$ and (b) $\delta=-0.5$. The profiles of panels (a) and (b) at several values of $\Delta$ are shown by the curves in (c) and (d). Other parameters used are $g=0.01J$ and $d=2$.
In Figs. 7(c) and 7(d), we plot the profiles of Figs. 7(a) and 7(b) at different frequency detunings in the region $\Delta\in(J,2J)$, respectively. It is found that, when $\delta=\pm 0.5$ and $\Delta=1.15J$ (the red solid curves), the spectra around the minimal reflection can be approximated as the asymmetric Fano line shapes. Here, the complete transmission ($R_{2}=0$) appears at $\Delta_{k}=\Gamma_{e}\{\sin(k_{\Delta}d)-[1+\cos(k_{\Delta}d)]\tan(2k_{\Delta}d-\phi_{k_{\Delta}})\}$. Then we can rewrite the reflection amplitude Eq. (LABEL:r2) as the superposition of two Lorentz spectra $r_{2,\pm}^{\text{AABB}}=r_{2,+}^{\text{AABB}}+r_{2,-}^{\text{AABB}}$, where
@@
We can further demonstrate that under the condition $\Gamma_{\pm}\gg\Gamma_{\mp}$, the reflection spectra around $\Delta_{\mp}$ can be fitted by the Fano line shapes as described by Eq. (40). According to Eq. (44b), we find that when the dimerization parameters $\delta=0.5$ ($-0.5$) and the frequency detuning $\Delta$ is taken as $\Delta/J\in(1.14,1.24)\cup(1.24,1.34)\cup(1.84,1.90)\cup(1.90,1.94)$ [$(1,1.01)\cup(1.31,1.39)\cup(1.39,1.47)\cup(1.87,1.92)\cup(1.92,1.95)$], we have $\Gamma_{+}\gg\Gamma_{-}$, and then the reflection spectra exhibit the Fano line shapes around the minimal reflection, as shown by the red solid curve in Fig. 7(c). When the dimerization parameters $\delta=0.5$ ($-0.5$) and $\Delta/J\in(1,1.02)\cup(1.52,1.58)\cup(1.58,1.71)\cup(1.99,2)$ [$(1.06,1.11)\cup(1.11,1.18)\cup(1.62,1.70)\cup(1.70,1.77)\cup(1.99,2)$], the width satisfy $\Gamma_{-}\gg\Gamma_{+}$. In this case, the reflection spectra exhibit the Fano line shapes around the minimal reflection, as shown by the purple dot-dashed curve in Fig. 7(c) [the red solid curve in Fig. 7(d)]. Here, the sign of $\delta$ affects the regions of the Fano line shapes, which is different from both the AAAA- and ABAB-coupling configurations.
When $\delta=0.5$ ($-0.5$) and $\Delta\approx 1.24J$ (the blue dashed curves), the reflection spectrum show the Lorentzian line shape (the super-Gaussian line shape). Specifically, for the Lorentzian line shape ($\delta=0.5$ and $\Delta\approx 1.24J$), we have the characteristic parameters $\Delta_{\text{S}L,k_{\Delta}}=\Delta_{\text{A}L,k_{\Delta}}\approx 0.99\Gamma_{e}$, $\Gamma_{\text{S},k_{\Delta}}\approx 3.38\Gamma_{e}$, $\Gamma_{\text{A},k_{\Delta}}\approx 0$, and $g_{\text{SA},k_{\Delta}}=\Gamma_{\text{SA},k_{\Delta}}=0$. In this case, the antisymmetric (symmetric) state is decoupled from (coupled to) the SSH waveguide. The reflection spectrum exhibits the Lorentzian line shape centered at $\Delta_{k}=0.99\Gamma_{e}$ with the linewidth of $3.38\Gamma_{e}$, as shown by the blue dashed curve in Fig. 7(c). For the super-Gaussian line shape, the reflection coefficient around the reflection peaks $\Delta_{k}=\Delta_{L2,k_{\Delta}}=\Delta_{L1,k_{\Delta}}=\Delta_{L,k_{\Delta}}$ can be approximated as
@@
where $\Delta^{\prime}=\Delta_{k}-\Delta_{L,k_{\Delta}}$. Equation (45) indicates that the reflection spectrum around the reflection peak exhibits a super-Gaussian line shape, as shown by the blue dashed curve in Fig. 7(d). When $\delta=\pm 0.5$ and $\Delta\approx 1.58J$ (the black dotted curves), the characteristic quantities become $\Delta_{\text{S}L,k_{\Delta}}=\Delta_{\text{A}L,k_{\Delta}}=\Gamma_{\text{S},k_{\Delta}}=\Gamma_{\text{A},k_{\Delta}}=g_{\text{SA},k_{\Delta}}=\Gamma_{\text{SA},k_{\Delta}}=0$. Therefore, both the symmetric and antisymmetric states are decoupled from the waveguide, and then the single photon is completely transmitted. This feature is consistent with the AAAA-coupling configuration. When $\delta=\pm 0.5$ and $\Delta\approx 1.70J$ (the purple dot-dashed curves), the reflection spectra show the asymmetric Fano ($\Gamma_{-}\gg\Gamma_{+}$) and Lorentzian line shape, respectively. For the Lorentzian line shape ($\delta=-0.5$ and $\Delta\approx 1.70J$), we have the characteristic parameters $\Delta_{\text{S}L,k_{\Delta}}=\Delta_{\text{A}L,k_{\Delta}}\approx-0.49\Gamma_{e}$, $\Gamma_{\text{S},k_{\Delta}}\approx 0$, $\Gamma_{\text{A},k_{\Delta}}\approx 0.51\Gamma_{e}$, $g_{\text{SA},k_{\Delta}}=\Gamma_{\text{SA},k_{\Delta}}=0$. In this case, the symmetric (antisymmetric) state is decoupled from (coupled to) the SSH waveguide. Therefore, the reflection spectrum exhibits the Lorentzian line shape centered at $\Delta_{k}=-0.49\Gamma_{e}$ with the linewidth of $0.51\Gamma_{e}$, as shown by the purple dot-dashed curve in Fig. 7(d).

### v2

@@ III.3.1 The AAAA-coupling configuration
[…]
In Fig. 5(b), we plot the profile of Fig. 5(a) at several frequency detunings in the region of $\Delta\in(J,2J)$; we find that when $\Delta\approx 1.20J$ (the blue dashed curve), the characteristic quantities in Eqs. (37a)–(37e) become $\Delta_{\text{S}L,k_{\Delta}}=\Delta_{\text{A}L,k_{\Delta}}=\Gamma_{e}$, $\Gamma_{\text{S},k_{\Delta}}=0$, $\Gamma_{\text{A},k_{\Delta}}=4\Gamma_{e}$, and $g_{\text{SA},k_{\Delta}}=\Gamma_{\text{SA},k_{\Delta}}=0$. Therefore, the symmetric (antisymmetric) state is decoupled from (coupled to) the waveguide. This feature results in the appearance of the Lorentzian line shape centered at $\Delta_{k}=\Gamma_{e}$ with the linewidth of $4\Gamma_{e}$. When $\Delta\approx 1.58J$ (the purple dot-dashed curve), the six characteristic quantities are $\Delta_{\text{S}L,k_{\Delta}}=\Delta_{\text{A}L,k_{\Delta}}=\Gamma_{\text{S},k_{\Delta}}=\Gamma_{\text{A},k_{\Delta}}=g_{\text{SA},k_{\Delta}}=\Gamma_{\text{SA},k_{\Delta}}=0$. In this case, both the symmetric and antisymmetric states are decoupled from the SSH waveguide. Therefore, the incident photon in the waveguide can be completely transmitted. When $\Delta=1.15J$ (the red solid curve) and $1.49J$ (the black dotted curve), the spectra become asymmetric, with a minimal reflection located at $\Delta_{k}=\Gamma_{e}\{\sin(k_{\Delta}d)-[1+\cos(k_{\Delta}d)]\tan(2k_{\Delta}d)\}$.
In the following analyses, we can show that in some regions the reflection spectrum around the minimal reflection can be approximated as a standard Fano line shape Fano1961 ; Miroshnichenko2010 . To this end, we decompose the reflection amplitude in Eq. (35b) as $r_{2,\pm}^{\text{AAAA}}=r_{2,+}^{\text{AAAA}}+r_{2,-}^{\text{AAAA}}$, where
@@
Figure 6: Reflection coefficient $R_{2}$ for the ABAB-coupling configuration as a function of $\Delta_{k}$ and $\Delta$. The dimerization parameter is (a) $\delta=0.5$ and (b) $\delta=-0.5$. The profiles of panels (a) and (b) at several values of $\Delta$ are shown by the curves in (c) and (d). Other parameters used are $g=0.01J$ and $d=2$.
@@ III.3.2 The ABAB-coupling configuration
For the ABAB-coupling configuration, the dimensionless parameters become $\mu_{1(3)}=\nu_{2(4)}=1$ and $\nu_{1(3)}=\mu_{2(4)}=0$ in Eqs. (35a)–(36f). To see the influence of $\Delta$ and $\delta$ on the single-photon scattering process, in Figs. 6(a) and 6(b) we show the reflection coefficient as a function of $\Delta_{k}$ and $\Delta$. Since the characteristic quantities contain the topology-dependent phase, it can be seen that the line shapes of the reflection spectra are sensitive to the sign of $\delta$, which is different from the AAAA-coupling configuration. To show the details more clearly, in Figs. 6(c) and 6(d), we plot the profiles of Figs. 6(a) and 6(b) at different frequency detunings in the region $\Delta\in(J,2J)$, respectively.
@@
We can see from the red solid curve in Fig. 6(c) [Fig. 6(d)] that, when $\delta=0.5$ ($-0.5$) and $\Delta\approx 1.20J$, the characteristic quantities are $\Delta_{\text{S}L,k_{\Delta}}=\Delta_{\text{A}L,k_{\Delta}}\approx-0.96\Gamma_{e}$ ($0.47\Gamma_{e}$), $\Gamma_{\text{S},k_{\Delta}}\approx 0$, $\Gamma_{\text{A}k_{\Delta}}\approx 2.82\Gamma_{e}$ ($0.46\Gamma_{e}$), and $g_{\text{SA},k_{\Delta}}=\Gamma_{\text{SA},k_{\Delta}}=0$. Therefore, the symmetric (antisymmetric) state is decoupled from (coupled to) the SSH waveguide. This makes the reflection spectrum exhibit the Lorentzian line shape centered at $\Delta_{k}=-0.96\Gamma_{e}$ ($0.47\Gamma_{e}$) with the linewidth of $2.82\Gamma_{e}$ ($0.46\Gamma_{e}$), as shown by the red solid curve. However, when $\delta=0.5$ ($-0.5$) and $\Delta\approx 1.58J$, i.e., the purple dot-dashed curve, the characteristic quantities become $\Delta_{\text{S}L,k_{\Delta}}=\Delta_{\text{A}L,k_{\Delta}}\approx-0.32\Gamma_{e}$ ($-0.95\Gamma_{e}$), $\Gamma_{\text{S},k_{\Delta}}\approx 7.79\Gamma_{e}$ ($5.26\Gamma_{e}$), $\Gamma_{\text{A}k_{\Delta}}\approx 0$, and $g_{\text{SA},k_{\Delta}}=\Gamma_{\text{SA},k_{\Delta}}=0$. In this case, the antisymmetric (symmetric) state is decoupled from (coupled to) the SSH waveguide, and then the reflection spectrum also exhibits the Lorentzian line shape but centered at $\Delta_{k}=-0.32\Gamma_{e}$ ($-0.95\Gamma_{e}$) with the linewidth of $7.79\Gamma_{e}$ ($5.26\Gamma_{e}$), as shown by the purple dot-dashed curve. When $\delta=\pm 0.5$ and $\Delta=1.49J$ (the black dotted curves), the spectra around the minimal reflection can be approximated as the asymmetric Fano line shapes. Here, the reflection spectrum takes the minima value ($R_{2}=0$) at $\Delta_{k}=\Gamma_{e}\{\sin(k_{\Delta}d-\phi_{k_{\Delta}})-[1+\cos(k_{\Delta}d-\phi_{k_{\Delta}})]\tan(2k_{\Delta}d)\}$. Specially, when $\delta=\pm 0.5$ and $\Delta\approx 1.27J$ (the blue dashed curves), the reflection spectra show the asymmetric Fano line shape and complete transmission, respectively. Similar to the AAAA-coupling configuration, we rewrite the reflection amplitude given by Eq. (35b) as the superposition of two Lorentz spectra $r_{2,\pm}^{\text{ABAB}}=r_{2,+}^{\text{ABAB}}+r_{2,-}^{\text{ABAB}}$, where the Lorentz-type amplitudes are given by
@@
Figure 7: Reflection coefficient $R_{2}$ for the AABB-coupling configuration as a function of $\Delta_{k}$ and $\Delta$. The dimerization parameter is (a) $\delta=0.5$ and (b) $\delta=-0.5$. The profiles of panels (a) and (b) at several values of $\Delta$ are shown by the curves in (c) and (d). Other parameters used are $g=0.01J$ and $d=2$.
To obtain the Fano line shapes for the reflection spectra, we need to ensure the condition $\Gamma_{\pm}\gg\Gamma_{\mp}$. Under this condition, the reflection spectrum near $\Delta_{\mp}$ can be approximated as the Fano line shape characterized by Eq. (40). Specifically, when the dimerization parameters $\delta=\pm 0.5$ and $\Delta$ is taken as $\Delta/J\in(1,1.02)\cup(1.48,1.58)\cup(1.58,1.68)\cup(1.99,2)$, the width $\Gamma_{\pm}$ satisfies $\Gamma_{+}\gg\Gamma_{-}$. Thus, the reflection spectra exhibit the Fano line shapes around the reflection minima, as shown by the black dotted curve in Figs. 6(c) and 6(d). When the dimerization parameters $\delta=\pm 0.5$ and $\Delta/J\in(1.11,1.20)\cup(1.20,1.27)\cup(1.27,1.30)\cup(1.82,1.89)\cup(1.89,1.94)$, we have $\Gamma_{-}\gg\Gamma_{+}$, and then the reflection spectrum exhibits the Fano line shape, as shown by the blue dashed curve in Fig. 6(c) (i.e., $\delta=0.5$ and $\Delta\approx 1.27J$). When $\delta=-0.5$ and $\Delta\approx 1.27J$, the characteristic quantities in Eqs. (37a)–(37e) become $\Delta_{\text{S}L,k_{\Delta}}=\Delta_{\text{A}L,k_{\Delta}}=\Gamma_{\text{S},k_{\Delta}}=\Gamma_{\text{A},k_{\Delta}}=g_{\text{SA},k_{\Delta}}=\Gamma_{\text{SA},k_{\Delta}}=0$. Therefore, both the symmetric and antisymmetric states are decoupled from the SSH waveguide. In this case, the input photon can be completely transmitted by the waveguide, as shown by the blue dashed curve in Fig. 6(d).
@@ III.3.3 The AABB-coupling configuration
For the AABB-coupling configuration, the dimensionless parameters in Eqs. (35a)–(36f) are $\mu_{1(2)}=\nu_{3(4)}=1$ and $\nu_{1(2)}=\mu_{3(4)}=0$. Figures 7(a) and 7(b) show the reflection coefficient as a function of $\Delta_{k}$ and $\Delta$. We find that the sign of $\delta$ can affect the reflection spectra of the AABB-coupling configuration, but not affect the location of the reflection peak ($R_{2}=1$). This is because the Lamb shifts $\Delta_{L1,k_{\Delta}}=\Delta_{L2,k_{\Delta}}=\Gamma_{e}\sin(k_{\Delta}d)$ are independent of the sign of $\delta$. The feature is different from the ABAB-coupling configuration.
In Figs. 7(c) and 7(d), we plot the profiles of Figs. 7(a) and 7(b) at different frequency detunings in the region $\Delta\in(J,2J)$, respectively. It is found that, when $\delta=\pm 0.5$ and $\Delta=1.15J$ (the red solid curves), the spectra around the minimal reflection can be approximated as the asymmetric Fano line shapes. Here, the complete transmission ($R_{2}=0$) appears at $\Delta_{k}=\Gamma_{e}\{\sin(k_{\Delta}d)-[1+\cos(k_{\Delta}d)]\tan(2k_{\Delta}d-\phi_{k_{\Delta}})\}$. Then we can rewrite the reflection amplitude Eq. (35b) as the superposition of two Lorentz spectra $r_{2,\pm}^{\text{AABB}}=r_{2,+}^{\text{AABB}}+r_{2,-}^{\text{AABB}}$, where
@@
We can further demonstrate that, under the condition $\Gamma_{\pm}\gg\Gamma_{\mp}$, the reflection spectra around $\Delta_{\mp}$ can be fitted by the Fano line shapes as described by Eq. (40). According to Eq. (44b), we find that when the dimerization parameters $\delta=0.5$ ($-0.5$) and the frequency detuning $\Delta$ is taken as $\Delta/J\in(1.14,1.24)\cup(1.24,1.34)\cup(1.84,1.90)\cup(1.90,1.94)$ [$(1,1.01)\cup(1.31,1.39)\cup(1.39,1.47)\cup(1.87,1.92)\cup(1.92,1.95)$], we have $\Gamma_{+}\gg\Gamma_{-}$, and then the reflection spectra exhibit the Fano line shapes around the minimal reflection, as shown by the red solid curve in Fig. 7(c). When the dimerization parameters $\delta=0.5$ ($-0.5$) and $\Delta/J\in(1,1.02)\cup(1.52,1.58)\cup(1.58,1.71)\cup(1.99,2)$ [$(1.06,1.11)\cup(1.11,1.18)\cup(1.62,1.70)\cup(1.70,1.77)\cup(1.99,2)$], the width satisfies $\Gamma_{-}\gg\Gamma_{+}$. In this case, the reflection spectra exhibit the Fano line shapes around the minimal reflection, as shown by the purple dot-dashed curve in Fig. 7(c) [the red solid curve in Fig. 7(d)]. Here, the sign of $\delta$ affects the regions of the Fano line shapes, which is different from both the AAAA- and ABAB-coupling configurations.
Figure 8: Reflection coefficient $R_{2}$ for the ABBA-coupling configuration as a function of $\Delta_{k}$ and $\Delta$. The dimerization parameter is (a) $\delta=0.5$ and (b) $\delta=-0.5$. The profiles of panels (a) and (b) at several values of $\Delta$ are shown by the curves in (c) and (d). Other parameters used are $g=0.01J$ and $d=2$.
When $\delta=0.5$ ($-0.5$) and $\Delta\approx 1.24J$ (the blue dashed curves), the reflection spectrum shows the Lorentzian line shape (the super-Gaussian line shape). Specifically, for the Lorentzian line shape ($\delta=0.5$ and $\Delta\approx 1.24J$), we have the characteristic parameters $\Delta_{\text{S}L,k_{\Delta}}=\Delta_{\text{A}L,k_{\Delta}}\approx 0.99\Gamma_{e}$, $\Gamma_{\text{S},k_{\Delta}}\approx 3.38\Gamma_{e}$, $\Gamma_{\text{A},k_{\Delta}}\approx 0$, and $g_{\text{SA},k_{\Delta}}=\Gamma_{\text{SA},k_{\Delta}}=0$. In this case, the antisymmetric (symmetric) state is decoupled from (coupled to) the SSH waveguide. The reflection spectrum exhibits the Lorentzian line shape centered at $\Delta_{k}=0.99\Gamma_{e}$ with the linewidth of $3.38\Gamma_{e}$, as shown by the blue dashed curve in Fig. 7(c). For the super-Gaussian line shape, the reflection coefficient around the reflection peaks $\Delta_{k}=\Delta_{L2,k_{\Delta}}=\Delta_{L1,k_{\Delta}}=\Delta_{L,k_{\Delta}}$ can be approximated as
@@
where $\Delta^{\prime}=\Delta_{k}-\Delta_{L,k_{\Delta}}$. Equation (45) indicates that the reflection spectrum around the reflection peak exhibits a super-Gaussian line shape, as shown by the blue dashed curve in Fig. 7(d). When $\delta=\pm 0.5$ and $\Delta\approx 1.58J$ (the black dotted curves), the characteristic quantities become $\Delta_{\text{S}L,k_{\Delta}}=\Delta_{\text{A}L,k_{\Delta}}=\Gamma_{\text{S},k_{\Delta}}=\Gamma_{\text{A},k_{\Delta}}=g_{\text{SA},k_{\Delta}}=\Gamma_{\text{SA},k_{\Delta}}=0$. Therefore, both the symmetric and antisymmetric states are decoupled from the waveguide and then the single photon is completely transmitted. This feature is consistent with the AAAA-coupling configuration. When $\delta=\pm 0.5$ and $\Delta\approx 1.70J$ (the purple dot-dashed curves), the reflection spectra show the asymmetric Fano ($\Gamma_{-}\gg\Gamma_{+}$) and Lorentzian line shape, respectively. For the Lorentzian line shape ($\delta=-0.5$ and $\Delta\approx 1.70J$), we have the characteristic parameters $\Delta_{\text{S}L,k_{\Delta}}=\Delta_{\text{A}L,k_{\Delta}}\approx-0.49\Gamma_{e}$, $\Gamma_{\text{S},k_{\Delta}}\approx 0$, $\Gamma_{\text{A},k_{\Delta}}\approx 0.51\Gamma_{e}$, and $g_{\text{SA},k_{\Delta}}=\Gamma_{\text{SA},k_{\Delta}}=0$. In this case, the symmetric (antisymmetric) state is decoupled from (coupled to) the SSH waveguide. Therefore, the reflection spectrum exhibits the Lorentzian line shape centered at $\Delta_{k}=-0.49\Gamma_{e}$ with the linewidth of $0.51\Gamma_{e}$, as shown by the purple dot-dashed curve in Fig. 7(d).